The perimeter of a square garden is 444m .what is its area

Answer by jim_thompson5910(35256)

P = 4s

444 = 4s

444/4 = s

111 = s

s = 111

-------------------------------------------------------

The side length is 111 meters.


A = s^2

A = 111^2

A = 12,321


The area is 12,321 square meters

------------------------------------------------------------------------

Let me know if that helps or not. Thanks.

If you need more help, feel free to email me at jim_thompson5910@hotmail.com

My Website: http://www.freewebs.com/jimthompson5910/home.html